this is a puzzle.we must fill in the empty spaces so that all the rows are true number sentences.i can't figure it out 4 1/2 ? 2 ? 1/3 =3 ,we must use order of operation
Answer by checkley71(8403)

(4*1)/2/2*3=3
(4/2)/2+3=3
(2/2)*3=3
1*3=3
3=3